OpenStudy (anonymous):

Calculate: (-2.2X10 -3)(-2.0X10 -4) Write your answer in scientific notation.

OpenStudy (unklerhaukus):

\[ (-2.2\times10^{ -3})(-2.0\times10 ^{-4})\\ =(-2.2\times-2.0)(10^{-3}\times10^{-4})\\ =(-2.2\times-2.0)(10^{-3-4})\\ =\]

OpenStudy (anonymous):

Calculate: \[(-2.2X10^{-3})(-2.0X10^{-4}) Write your answer in scientific notation.

OpenStudy (anonymous):

So 4.4 X 10\[^{-7}\] is this right?

OpenStudy (unklerhaukus):

yes

OpenStudy (anonymous):

Awesome! Thanks!